OpenStudy (anonymous):

will fan and medal What is the area of this rectangle in square feet? A. 72 square feet B. 216 square feet C. 24 square feet D. 288 square feet

OpenStudy (anonymous):



OpenStudy (igreen):

The formula for the area of a rectangle is \(l \times w\). So what's \(4 \times 6\)?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

but whats hard is i need it in square inches

OpenStudy (anonymous):

i mean square feet

OpenStudy (igreen):

Oh, nevermind..we have to change the measurements.

OpenStudy (anonymous):

yep

OpenStudy (igreen):

1 yard is 3 feet. So just multiply: 4 * 3 = ? 6 * 3 = ?

OpenStudy (anonymous):

4*3=12 6*3=18

OpenStudy (igreen):

Yep, so those are your measurements.. Now multiply \(\sf 12ft \times 18ft\), what do you get?

OpenStudy (anonymous):

216

OpenStudy (anonymous):

so its 216^2ft right

OpenStudy (anonymous):

^2 meaning squared

OpenStudy (igreen):

Yep!

OpenStudy (anonymous):

i will let u know my score k

OpenStudy (igreen):

\(\sf 12ft \times 18ft = 216ft^2~ \Large\color{lime} \checkmark\)

OpenStudy (anonymous):

75 due to the fact that i did most of it myself

OpenStudy (anonymous):

as in previouse questions

OpenStudy (anonymous):

thx